### v1

# Probing Politico-Economic Bias in Multilingual Large Language Models: A Cultural Analysis of Low-Resource Pakistani Languages

Abstract
Large Language Models (LLMs) are increasingly shaping public discourse, yet their politico-economic biases remain underexamined in non-Western and low-resource multilingual contexts. This paper presents a systematic analysis of political bias in 13 state-of-the-art LLMs across five low-resource languages spoken in Pakistan—Urdu, Punjabi, Sindhi, Balochi, and Pashto. We propose a novel framework that integrates an adapted Political Compass Test (PCT) with a multi-level framing analysis. Our method combines quantitative assessment of political orientation across economic (left–right) and social (libertarian–authoritarian) axes with qualitative analysis of framing through content, style, and emphasis. We further contextualize this analysis by aligning prompts with 11 key socio-political themes relevant to Pakistani society. Our results reveal that LLMs predominantly align with liberal-left values, echoing Western training data influences, but exhibit notable shifts toward authoritarian framing in regional languages, suggesting strong cultural modulation effects. We also identify consistent model-specific "bias signatures" and language-conditioned variations in ideological expression. These findings show the urgent need for culturally grounded, multilingual bias auditing frameworks.
Probing Politico-Economic Bias in Multilingual Large Language Models:
A Cultural Analysis of Low-Resource Pakistani Languages
Afrozah Nadeem, Mark Dras, Usman Naseem School of Computing, Macquarie University, Australia, afrozah.nadeem@students.mq.edu.au, {mark.dras,usman.naseem}@mq.edu.au
1 Introduction
Large Language Models (LLMs) have rapidly transformed the field of Natural Language Processing (NLP), demonstrating impressive capabilities across a wide array of linguistic and cultural contexts Blodgett et al. (2020). As these models become increasingly integrated into daily life, growing concerns have emerged about their potential to propagate social harms, including hallucinations, gender, and racial biases Zheng et al. (2023). Among these concerns, political bias stands out as particularly consequential, given its potential to misrepresent public discourse, reinforce dominant ideologies, and exacerbate societal polarization Demszky et al. (2019).
LLMs often reflect the socio-political assumptions embedded in their training data, frequently mirroring political biases aligned with high-resource, Western contexts Weidinger et al. (2021). While significant research has focused on political bias in models operating in English, Chinese, and other high-resource languages, relatively little attention has been paid to how such biases manifest in low-resource languages—especially in multilingual, culturally complex regions like South Asia Kumar et al. (2023). The intersection of linguistic diversity and socio-political dynamics in low-resource languages poses unique challenges for building fair and culturally attuned AI systems Suarez et al. (2019).
Pakistan presents a particularly compelling case for such investigation Harman (2018). With five major low-resource languages—Urdu, Punjabi, Sindhi, Balochi, and Pashto—spoken by over 80% of the population, Pakistan encompasses diverse political ideologies shaped by religion, regional history, and socio-economic disparities Sun et al. (2022). Sensitive topics such as blasphemy laws, minority rights, language policies, and provincial autonomy dominate political discourse, making it imperative to understand how LLMs engage with such issues across different linguistic and cultural dimensions.
Existing research on political bias in LLMs suffers from several important limitations Naveed et al. (2024). First, current studies largely focus on Western languages and political paradigms, leaving non-Western contexts critically underexplored Chen et al. (2020). Second, most work isolates stance detection from the broader question of how political orientations are linguistically framed in generated content Yu et al. (2023). Third, current evaluation frameworks rarely account for the cultural nuances and socio-political particularities that shape language use in South Asia Bordia and Bowman (2019). While some studies have examined multilingual LLMs, they often treat languages as independent variables rather than analyzing how political ideologies may transfer, transform, or intensify across culturally linked but linguistically distinct languages Bang et al. (2021).
This research addresses these gaps by introducing a comprehensive, culturally grounded framework for evaluating political and economic bias in multilingual LLMs, focusing specifically on five low-resource Pakistani languages. We propose a novel two-tier methodology that combines an adapted Political Compass Test (PCT) with a fine-grained framing analysis that examines both what is said (stance) and how it is said (style), as shown in Figure 1. Our approach captures ideological bias not only in the LLMs’ political positions but also in their narrative framing, across languages that reflect diverse cultural and political realities. This study makes the following key contributions:
  • We conduct the first large-scale evaluation of political bias in LLMs across five low-resource Pakistani languages.
  • We adapt and manually translate the PCT into these languages and align it with 11 culturally salient topics such as LGBTQ+ rights, blasphemy laws, and language policy, grounded in the Pakistani socio-political context.
  • We introduce a three-pronged framing analysis that integrates Boydstun’s frame dimensions, NER, and lexical polarity assessment to dissect both content and stylistic framing biases.
  • We analyze 13 SOTA LLMs to investigate how political positions and framing strategies vary across languages, and how linguistic choice activates culturally specific ideological shifts.
2 Related Work
Political Bias in Language Models: The political orientation of LLMs has become a core issue in AI ethics. Early work by Liu et al. (2021) showed that word embeddings reflect human-like ideological bias. Later studies, such as Motoki et al. (2024) and Ceron et al. (2024), demonstrated that models like GPT-3/4 tend to lean liberal on social matters and can convincingly emulate partisan ideologies. However, these findings are largely based on U.S.-centric frameworks, limiting their cross-cultural relevance.
Evaluation Tools: Political Compass and Policy Probes: The Political Compass Test (PCT) has been widely used for evaluating AI political alignment Hartmann et al. (2023), offering a structured assessment across ideological axes. Others like Bang et al. (2021) introduced alternative methods based on policy preference tasks. Yet, these frameworks are developed for English-speaking contexts, often assuming direct translatability across languages and cultures—an assumption that fails in multilingual and low-resource settings.
Bias in Multilingual and Low-Resource Contexts: Research has shown that traditional bias detection tools underperform in non-English environments due to linguistic and cultural misalignment Barkhordar et al. (2024). South Asian languages such as Punjabi, Sindhi, Pashto, and Balochi are especially underserved in LLM training data. In other study Harman (2018) further found that sentiment models trained on English data degrade when applied to translated content, underscoring the need for culturally-aware evaluation strategies. Recent work also cautions against treating multilingual bias as language-independent Thapa et al. (2024).
Framing and Discourse-Level Analysis: Most political bias research emphasizes stance detection, neglecting how bias manifests through rhetorical framing. The taxonomy by Boydstun et al. (2014) provides a foundation for deeper analysis of issue framing, yet remains underused in LLM evaluations. Framing is particularly relevant in multicultural settings, where political language varies not just in content but in style, tone, and structure—areas that remain largely unexplored in multilingual NLP.
Figure 1: Overview of our proposed framework for political bias analysis for evaluating political bias in language models. The framework features a political compass approach for stance detection and decomposes bias into content and style dimensions, examining controversial topics across Pakistani languages.
Despite recent progress, several critical limitations remain in the study of political bias in language models. Much of the existing literature is Western-centric, focusing predominantly on high-resource languages like English and Spanish, with limited attention to multilingual or non-Western contexts. Moreover, while stance detection is a common approach, it often overlooks how political bias is conveyed through rhetorical framing. Tools like the PCT, though widely used, have not been culturally adapted for regions like South Asia, limiting their interpretability and relevance. In response to these gaps, our study makes several key contributions. We culturally adapt the PCT for five low-resource Pakistani languages—Urdu, Punjabi, Sindhi, Balochi, and Pashto—ensuring linguistic accuracy and ideological relevance. We propose a novel three-layered framing analysis that integrates policy frame taxonomy, named entity recognition (NER), and lexical polarity scoring to capture both content and style. Our evaluation spans 11 culturally salient political topics central to Pakistani discourse, and we benchmark 13 leading language models, both open-source and commercial. To our knowledge, this is the first comprehensive investigation of political bias in LLMs conducted within a multilingual, non-Western framework.
3 Method
Overview: Figure 1 presents the comprehensive methodological framework underpinning our political bias analysis. First, we focus on the quantitative evaluation of political stance, utilizing the Political Compass Test to position model outputs across economic and social dimensions. Second, we expand the analysis by examining framing bias, incorporating content framing, named entity recognition, and lexical polarity to capture the nuanced ways models express ideological positions within culturally relevant political topics. Each tier is discussed in detail in the following sections.
3.1 Political Leaning of pretrained LLM
Overview and Motivation: To evaluate political bias in large language models (LLMs), we adopt a two-step evaluation approach grounded in political spectrum theories, following the framework proposed by Feng et al. (2023). This framework characterizes political opinion using a two-axis model: social values, ranging from liberal to conservative ($S_{\mathrm{soc}}$), and economic values, ranging from left to right ($S_{\mathrm{eco}}$). This dual-dimensional representation offers a nuanced view of political ideologies, surpassing the simplistic left-right dichotomy and enabling deeper analysis of how political bias manifests in LLMs Liu et al. (2024).
Political Compass Test (PCT): The PCT [^1] https://www.politicalcompass.org/test consists of 62 statements, denoted as $P_{\mathrm{s}}$, designed to assess a respondent’s political orientation across the $S_{\mathrm{eco}}$ and $S_{\mathrm{soc}}$ axes Smith et al. (2022). Each participant evaluates the statements using a four-level Likert scale: Strongly Agree, Agree, Disagree, and Strongly Disagree. The final scores are calculated based on weighted responses and mapped to a coordinate in the range $[-10,10]$ for each axis, producing a two-dimensional representation of political ideology.
Translation and Human Annotation: To adapt the PCT for low-resource languages in Pakistan, we employed native speakers with expertise in political discourse from five different cities. Each statement was translated into five Pakistani languages, detailed Table in Appendix 18: Urdu– national language, Punjabi – most widely spoken regional language, Sindhi– western provincial language, Balochi – southwestern provincial language,Pashto – northwestern provincial language.
Language Models and Prompt Design: We evaluated thirteen LLMs categorized into open-source and closed-source models. The closed-source models include: OpenAI, OpenAI o1-mini, OpenAI o1-preview, GPT-3.5-turbo, GPT-4, GPT-4-turbo, GPT-4o [^2] https://openai.com/api, Claude [^3] https://www.anthropic.com, and Gemini 1.5 Pro [^4] https://ai.google.dev/gemini-api/docs. The open-source models include: Mistral-7B-Instruct [^5] https://docs.mistral.ai/api/, DeepSeek [^6] https://lambda.ai/, RoBERTa-large [^7] https://huggingface.co/FacebookAI/xlm-roberta-large, RoBERTa-base11, and BERT-large 12. For each language, we developed tailored prompts to generate high-quality responses, which are shown in Appendix 15.
 [^8] https://huggingface.co/FacebookAI/xlm-roberta-base
Response Generation and Parameters: To ensure consistency, we fixed generation parameters across all models: $\text{Temperature}=0.5$, Top-$p$ = 1. All responses were generated in a zero-shot multilingual setting. For stance detection, we used the multilingual mDeBERTa-v3-base model, which supports over 100 languages, including all five under study.
Stance Scoring and Mapping: To quantify the model’s political stance on each PCT statement, we employ a two-stage computational approach. First, we calculate a scalar stance score from classifier confidence values. Second, we discretize this continuous score into ordinal categories that align with the PCT response format. This allows us to convert raw classifier outputs into interpretable symbolic judgments.
Let $A_{s},A,D,D_{s}\in[0,1]$ denote the predicted confidence scores corresponding to the classes Strongly Agree, Agree, Disagree, and Strongly Disagree, respectively.
1. Stance Score Computation: We define a scoring function $f:[0,1]^{4}\rightarrow[-10,10]$ to assign a value based on the dominant confidence class:
$f(A_{s},A,D,D_{s})=\begin{cases}10\cdot A_{s}&\text{if }A_{s}>\max(A,D,D_{s})\\ 5\cdot A&\text{if }A>\max(A_{s},D,D_{s})\\ -10\cdot D_{s}&\text{if }D_{s}>\max(D,A,A_{s})\\ -5\cdot D&\text{otherwise}\end{cases}$
The result $S=f(A_{s},A,D,D_{s})\in[-10,10]$ serves as a continuous stance score, indicating both the direction and strength of agreement.
2. Stance Discretization: To facilitate comparative analysis across models and statements, we define a discretization function $g:[-10,10]\rightarrow\{0,1,2,3\}$ that maps the continuous score to categorical labels using a symmetric threshold parameter $\tau>0$:
$g(S)=\begin{cases}3&\text{if }S\geq 2\tau\\ 2&\text{if }0\leq S<2\tau\\ 1&\text{if }-2\tau<S<0\\ 0&\text{if }S\leq-2\tau\end{cases}$
This results in an ordinal stance label interpreted as:
  • 3 – Strongly Agree
  • 2 – Agree
  • 1 – Disagree
  • 0 – Strongly Disagree
By mapping soft classifier outputs to these well-defined categories, we ensure that downstream aggregation and political leaning visualization remain interpretable and robust. This method also permits consistency across languages and LLMs in our multilingual evaluation setting. The final stance scores across all statements are aggregated for each model and projected onto the two-dimensional ($S_{\mathrm{eco}}$, $S_{\mathrm{soc}}$) space. This facilitates a structured evaluation of political alignment and model behavior across both ideological dimensions.
3.2 Ideological Framing Analysis
While the PCT quantifies political orientation along economic and social dimensions, it lacks detailed insight into how these ideologies are expressed in discourse Rozado (2024). To address this, we propose an ideological framing analysis framework that examines how large language models (LLMs) communicate politically sensitive topics through content and stylistic choices Liu et al. (2024). This method complements PCT by analyzing not only the stance but also the narrative strategies LLMs employ to present their positions Abdurahman et al. (2024).
Topic Selection and Data Generation: We select eleven politically salient topics in the Pakistani context, each mapped to PCT dimensions and characterized by polarized opinions. Topics were identified based on prior research Lee et al. (2022), reputable institutions (e.g., Pew Research Center [^9] https://www.pewresearch.org/topics/), and media bias trackers (e.g., Allsides.com [^10] https://www.allsides.com/topics-issues) Bang et al. (2024). The topics, denoted $T=\{T_{(i)}\}$, include:
LGBTQ Rights and Same-Sex Marriage, Blasphemy Laws, Education, Freedom of Press, Abortion Rights, Death Penalty, Climate Change Ejaz et al. (2023), Language Policy, Welfare and Charity, Religious Minorities Rights, Policing and Surveillance.
Each topic was translated into five low-resource Pakistani languages to enable multilingual evaluation.
For each topic $T_{(i)}$, we generate news headlines $H_{\mathrm{LLM}}$ in two opposing stances: proponent $T_{(i)\mathrm{pro}}$ and opponent $T_{(i)\mathrm{opp}}$. Headlines are an ideal unit for framing analysis as they encapsulate the core message and tone of discourse Lee et al. (2022); Sheng et al. (2021); Baly et al. (2020). We generated 1000 headlines per stance, per language, using prompts that explicitly specify stance to elicit contrasting viewpoints Nadeem et al. (2021) (see Appendix  15 for prompt details).
Frame Dimension Classification: To investigate the ideological narratives, we classify generated headlines using Boydstun’s 15 cross-cutting frame dimensions Boydstun et al. (2014), which capture salient political discourse themes such as “Economics,” “Morality,” “Health and Safety,” and “Cultural Identity.” These topic-agnostic frames facilitate comparative analysis across models and topics Hamborg (2020), for details, see Appendix D. Frame classification was performed using GPT-3.5-turbo with bilingual prompts, computing frame ratios to identify dominant framing strategies employed by each model Ziems and Yang (2021).
Entity-Based Framing Analysis: Beyond content frames, we analyze framing at the entity level by extracting and quantifying frequently mentioned named entities (e.g., political figures, countries, organizations) from generated headlines. Frequent entity mentions serve as a subtle framing device, indicating which actors or groups models emphasize Devlin et al. (2019); Zheng et al. (2023). NER was conducted using a pretrained bert-base-multilingual-cased model  [^11] https://huggingface.co/google-bert/bert-base-multilingual-cased.
Sentiment Polarity Towards Entities: To assess stylistic and attitudinal bias, we analyze the sentiment polarity expressed towards entities within headlines. Using a fine-tuned XLM-RoBERTa model for Urdu sentiment analysis Fan et al. (2019), we categorize sentiment as positive, negative, or neutral Saez-Trumper et al. (2013). Aggregating sentiment across samples yields a profile of the model’s overall stance tone towards each topic, revealing subtle stylistic bias Spliethöver et al. (2022); Roy and Goldwasser (2020).
4 Experimental Settings
Dataset: We introduce and release a novel multilingual dataset designed to support political bias analysis across five low-resource Pakistani languages: Urdu, Punjabi, Sindhi, Balochi, and Pashto [^12] https://anonymous.4open.science/r/PoliticalBiasEvaluation-10DE. The dataset comprises two main components:
  • Political Compass Test (PCT) Translations: This segment includes 62 culturally adapted political statements translated into each of the five target languages. Responses were collected from 13 large language models (LLMs), generating a total of 4,030 responses (62 statements × 13 models × 5 languages). All translations were produced and validated by native speakers, with triple verification to ensure full linguistic and semantic fidelity. Inter-annotator agreement for these translations achieved a Fleiss’ kappa score of 0.99, indicating near-perfect consensus.
  • Headline Generation Corpus: To evaluate framing bias, we generated news headlines using four SOTA LLMs across 11 politically sensitive topics in both proponent and opponent stances tags see detail in Appendix 2. Each model produced 22,000 headlines per language, resulting in a total of approximately 444,340 multilingual headlines (22,000 × 4 models × 5 languages). A stratified 20% sample (2,200 headlines) was manually evaluated for linguistic correctness and semantic coherence, yielding a 100% language correctness rate and a semantic agreement rate of 98.5%. The corresponding Fleiss’ kappa score for semantic consistency in the DeepSeek model was 0.34, reflecting moderate inter-annotator agreement given the subjective nature of generated content.
The dataset addresses a critical gap in non-Western language resources for political discourse analysis and AI bias evaluation. All annotators were recruited from linguistically representative regions—Lahore (Urdu), Sahiwal (Punjabi), Karachi (Sindhi), Quetta (Balochi), and Peshawar (Pashto)—and were compensated fairly.
Category Details
PCT
Political statements 62 $P(s)$
Languages 5
LLMs 13
$R_{\text{LLM}}$ 4,030
Agreement score Fleiss’ kappa = 0.99 (near-perfect)
Headline Generation Corpus
Political topics 11 $T(i)$
Stance tags $T(i)_{\text{Pro}}$, $T(i)_{\text{Opp}}$
LLMs 4
Headlines per language per model 22,000
$H_{\text{LLM}}$ 444,340 (approx.)
Human evaluation sample size 2,200 headlines (20% sample)
Language correctness rate 99%
Semantic agreement rate 98.5%
Table 1: Summary Statistics for Political Compass Test Translations and Headline Generation Corpus
Computational Resources: This study demanded significant computational resources to evaluate political bias through multiple language models and linguistic contexts. We utilized commercial API services for model access, with OpenAI API costs approximately \$287 USD for open-AI model evaluations (GPT-3.5-turbo, GPT-4, GPT-4-turbo, open-ai mini, open-ai preview) in five languages. Additional computational expenses included cloud computing resources for running open-source models and multilingual NLP pipelines for entity identification and sentiment analysis. The total computational budget, including API costs, cloud infrastructure, and model inference, was estimated at \$350 USD. Processing the complete experimental pipeline, including PCT evaluation, 440,340 headline generation, and three-method framing analysis, required approximately 120 GPU-hours on NVIDIA A100 instances.
5 Results and Analysis
5.1 Political Stance Distribution Across Languages
Figure 2 presents political stance detection results for various LLMs evaluated in Urdu, Punjabi, Balochi, Sindhi, and Pashto using the Political Compass Test (PCT). The majority of models cluster within the libertarian-left quadrant, indicating progressive economic views coupled with socially independent positions. Among them, Claude exhibits the strongest libertarian tendencies, while GPT-4-turbo scores the most pronounced leftward economic stance.
Interestingly, some models, such as OpenAI o1-mini and GPT-3.5-turbo, position in the authoritarian-right quadrant when responding in Urdu, indicating traditional or conservative leanings in this language context. Sindhi results largely reflect libertarian-left stances across models.
Overall, the PCT results reveal considerable variance in political positioning across models and languages. GPT-based models consistently shift toward authoritarian-left stances in regional languages (Urdu, Punjabi, Pashto, Balochi), whereas open-source models like Claude, Gemini, Mistral, and DeepSeek predominantly maintain libertarian-left orientations Johnson and Goldwasser (2016). This suggests that GPT models may interpret non-Western linguistic inputs as favoring stronger state control.
BERT variants generally display right-leaning tendencies irrespective of language. These findings underscore how language choice significantly influences LLMs’ expressed political biases, reinforcing the necessity for multilingual evaluations to fully capture bias dimensions.
Figure 2: Political leaning of open source and closed source models used for low-resource language shows diverse inclination across LLM
5.1.1 Cultural Adaptive Models are Less Biased
We have performed additional experiments on the Urdu language to evaluate the political bias on LLMs, to analyse it, we fine-tuned models: Bert-base [^13] https://huggingface.co/google-bert/bert-base-cased,Gemma7b [^14] https://huggingface.co/google/gemma-7b, GPT-3.5, and Roberta-small—all specifically adapted for the Urdu language as shown in Figure 3. The model political inclination can be verified by the bias score of a language model based on its political positioning in a 2D ideological space, which can be measured as:
Bias $\text{Bias Score}=\sqrt{(x-0)^{2}+(y-0)^{2}}=\sqrt{x^{2}+y^{2}}$
Where:
  • $x$ = position on the economic scale
  • $y$ = position on the social (authoritarian/libertarian) scale
  • $(0,0)$ = perfect neutrality (bias score = 0)
Other bias score formulations are Bias Score = |x| + |y|  [^15] Manhattan Distance or simple magnitude Bias Score = Max(|x|, |y|)  [^16] Chebyshev Distance. The interpretation scale is:
Figure 3: Political leaning of four open-source LLMs used for the Fine-tuning on Urdu language: a deep analysis on cultural nuance and political inclination of LLMs
$\textbf{LM}=\left\{\begin{array}[]{ll}\text{Perfect neutrality}&\text{if Bias Score }=0\\ \text{Highly neutral}&\text{if Bias Score }<1\\ \end{array}\right.$
Model classification by bias score is measured as:
LM =$\textbf{LM }=\left\{\begin{array}[]{ll}\text{Neutral model}&\text{if Bias Score }\in[0,1)\\ \text{Biased model}&\text{if Bias Score }\geq 1\\ \end{array}\right.$
The closer the bias score is to zero, the more politically neutral the model is. Any deviation from the origin $(0,0)$ indicates increasing political bias. This mathematical framework clearly shows that zero represents perfect neutrality, and any movement away from the center introduces measurable bias.
The Urdu fine-tuned models demonstrated superior performance in realizing more composed and centered political positioning compared to state-of-the-art (SOTA) models. These models effectively reduced political bias and achieved improved neutrality. This outcome indicates that targeted fine-tuning on Urdu linguistic features helps models to better engage with nuanced cultural and political discourse. These models show knowledgeable and context-aware behavior in terms of political and cultural perceptions rooted specifically in Pakistan. The major implication of Fine-tuning for cultural $\rightarrow$ linguistic adaptation can integrate regional political viewpoints, producing models that are culturally aware while remaining politically neutral. This highlights the necessary balance between cultural relevance and political neutrality in AI systems.
Figure 4: Frame dimension ratio for "Religious minority rights" topic for four models. Overall similar but most variance observed in "Morality", "Fairness and Equality" frames.
Figure 5: Entity-Based Frame Analysis on Pakistan’s Religious Rights. The left side shows performance comparison of LLMs (GPT-4, Gemini, DeepSeek, Claude) analyzing Pakistan discourse. The right side presents top 10 entities with sentiments showing Positive, Negative, Neutral. Supreme Court and National Commission as central nodes. Green circles like Interior Minister indicate positive sentiment, while red nodes including Parliament and Rights reflect critical discourse. Regional entities Punjab, Sindh, and Balochistan connect to the central Pakistan core. Dashed circles represent unique entities. This visualization maps institutional relationships shaping religious minority protections in Pakistan’s governance framework.
5.2 Framing Bias Analysis
5.2.1 Content Bias
Framing Dimension: All models employ cultural identity frames, but with varying concentrations, the distribution specifies how models differently interpret religious minority issues in Pakistan’s context, with Claude taking a universalist rights approach versus others incorporating more culturally-specific religious and social considerations. The cultural aspects reveal distinct patterns: considering along a specific topic, where Figure 4 illustrates the distribution of Boydstun’s frame dimensions for the Religious Minority Rights topic (opponent stance) in Urdu across four LLMs. Distinct cultural framing patterns emerge: Claude primarily employs fairness/equality frames, emphasizing rights-based discourse. In contrast, GPT-4 balances morality, fairness, and external regulation frames, suggesting a multifaceted interpretive lens. For the Death Penalty topic (opponent stance), cultural framing biases differ. Claude and Gemini prioritize morality frames, embedding religious and ethical arguments. GPT-4 and DeepSeek focus on policy evaluation and fairness frames. Across all models, constitutional and jurisprudence frames appear with similar frequency, reflecting Pakistan’s Islamic legal framework as a shared contextual baseline. Minimal use of cultural identity frames suggests models approach the death penalty predominantly via universal human rights discourse, with morality providing the primary ideological differentiation Fazal (2022).
Entity Frequency Analysis: Entity frequency analysis for Religious Minority Rights reveals framing patterns aligned with political orientations Schramowski et al. (2022). Figure  5 shows that models exhibiting more authoritarian stances tend to frame minority rights through institutional and geographic hierarchies, emphasizing state and regional across top 10 entities, such as Pakistan, the Supreme Court, and regional bodies serve to situate the discourse within legal and geographic contexts, words like rights and law reinforce a rights-based framing. Libertarian-leaning models present a broader spectrum of entities, incorporating both legal frameworks and regional minority experiences, indicating a more nuanced framing.
Figure 6: Deep Seek model topic by polarization in Urdu Headlines: Topics ranked by opinion extremity, with color dots indicating positive (green) or negative (red) sentiment direction.
5.2.2 Stylistic Bias
Lexical Polarity: Lexical polarity analysis highlights how LLMs stylistically frame within Pakistani political discourse. Figure 6 discloses captivating polarization patterns in Urdu media headlines through numerous topics, where Same sex marriage rights producing the most extreme coverage with positive sentiment, followed closely by language policy, welfare charity, and education topics. Whereas, religious and moral issues, abortion rights, the death penalty, and blasphemy laws show high polarization with negative sentiment. Government performance topics show moderate-high polarization with negative sentiment. The heatmap compares sentiment patterns between proponents and opponents of religious minorities’ rights, considering a topic specifically related to Religious Minority Rights. Figure 7 shows predominantly positive sentiment, while opponents display more polarized views. This indicates that LLMs frame religious minority rights using constructive and humanitarian language, consistent with international human rights norms, while recognizing Pakistan’s complex religious landscape. Such stylistic tendencies suggest a diplomatic approach by LLMs to sensitive issues, emphasizing rights and dialogue over conflict or aggression.
Figure 7: Heatmap comparing sentiment patterns between proponents and opponents of religious minorities’ rights. Proponents show predominantly positive sentiments, while opponents display more polarized views with negative sentiment alongside moderate distribution across other categories. Both groups maintain similar neutral sentiment levels.
5.3 Model-Level Bias Interpretation
To further interpret how political bias manifests in model behavior, we conducted a detailed, multidimensional analysis of the DeepSeek model’s outputs. This includes political compass positioning, stance-specific framing differences, sentiment polarity rates, and topic-wise alignment patterns. The analysis confirms that DeepSeek consistently occupies a left-libertarian space while framing proponent content more positively and emphasizing tradition in opponent discourse. A sequential flowchart outlines the model’s internal decision process across stance, framing, and polarity layers. More detailed visual and quantitative breakdowns are provided in the Appendix  14.
5.4 Cross-Cultural model insights
Our analysis discloses discrete model traits that interact differently with cultural-linguistic contexts. Cross-culturally, we identified the detailed experimental political appearance where cultural schematic initiation increased religious authority orientations, framing decomposition adaptation based on entity emphasis, and sentiment variation can reduce positivity in controversial topics. GPT models determine high cultural sensitivity, fluctuating from libertarian-left in English to authoritarian-left in Pakistani languages, with influential framing inclinations. On the other hand, Claude conserves stable libertarian- positioning across languages with higher progressive sentiment, demonstrating Western values. Open-source models show diverse patterns: Deep Seek exhibits moderate variation while maintaining dependable entity selection, and Gemini shows prominent ideological framing differences. This challenges "universal" bias mitigation approaches, suggesting that effective global AI deployment requires balancing universal principles with cultural sensitivity—a balance current LLMs struggle to achieve, exhibiting either rigid Western liberalism or potentially excessive cultural adaptation.
5.5 Multilingual Influence on Political Bias
To analyse the results of political stance detection plots for all languages, the clustering plots in different quadrants provide deep insight that the model itself behaves in a biased manner depending on political context, and shifts its opinions. This reflects an imbalance in the LLMs’ training data, highlighting that English-dominant models may develop different political contexts in lower-resource languages.
The models respond to culturally-specific political framing—particularly on topics like religious and minority rights—demonstrating language-specific implications. The most significant comparison is shown in Figure 13, where the models reveal both consistent patterns across topics, such as universal support for Education, and prominent divergence for Language policy.
This observation underscores how multilingualism can create language-specific political personalities within the same LLM architecture. Notably, this is a finding with thoughtful implications for future development, as global deployment may unintentionally reinforce different political biases depending on the language of interaction.
6 Conclusion
This study presents a comprehensive analysis of political bias in multilingual LLMs across five low-resource Pakistani languages. Using a two stage framework—an adapted Political Compass Test and multi-level framing analysis—we evaluated 13 LLMs and found that political stances vary systematically with language. While most models reflect liberal-left biases in English, regional languages often shift responses toward authoritarian or culturally specific framings. These findings reveal that language acts as a key cultural modulator of bias and that models exhibit consistent bias patterns across linguistic contexts. Our work highlights the importance of culturally grounded evaluation in assessing LLM behavior.
Limitations
This study faces numerous methodological and scope-related restraints that should be acknowledged. First, our analysis is restricted to five Pakistani languages, potentially missing important variations in other South Asian languages like Bengali, Tamil, or Persian that might reveal different bias patterns. The Political Compass Test, while widely used, was designed for Western political contexts and may not fully capture the nuances of Pakistani political dimensions, such as the role of Islamic jurisprudence or tribal governance structures. Moreover, our entity recognition and sentiment analysis tools were primarily trained on English data, potentially reducing their accuracy for Pakistani languages and missing culturally-specific entities or sentiment expressions. The temporal limitation of our training data means we cannot account for how recent political events might influence model behaviours. To developing bias mitigation techniques that preserve cultural sensitivity while reducing harmful stereotypes presents a crucial challenge, possibly through confrontational training methods or constitutional AI approaches adapted for multilingual contexts.
Ethical Statement
This research was conducted with strict devotion to ethical principles, ensuring cultural sensitivity and participant welfare. Content generation carefully avoided potentially harmful or inflammatory material while maintaining analytical integrity. We acknowledge possible biases in our Western-developed evaluation frameworks and commit to transparent reporting of limitations. The dataset excludes personally identifiable information and extreme political content that could incite violence or discrimination. We recognize the responsibility of AI bias research in multicultural contexts and emphasize that our findings should inform inclusive AI development rather than reinforce stereotypes. This work aims to promote reasonable AI systems that respect diverse political perspectives and cultural values across Global South communities.
Appendix A Appendix
Dataset Contribution Our research provides a valuable multilingual dataset that spans five Pakistani languages (Urdu, Punjabi, Sindhi, Balochi, and Pashto) that can serve as a basis for future political bias and linguistic studies. For the Political Compass Test, the statement are translated into five low-resource languages shown in the Table 18. To work with five Pakistani languages (Urdu, Pashto, Sindhi, Balochi, and Punjabi), the design study proposed three annotators from the specific region of Pakistan, and each of them are highly expert in speaking and writing in the dedicated low-resource language of their area. The dataset addresses a critical gap in non-Western language resources for political discourse analysis and AI bias evaluation. We chose the Method: Triple-verified by native speakers. All annotators were recruited from linguistically representative regions—Lahore (Urdu), Sahiwal (Punjabi), Karachi (Sindhi), Quetta (Balochi), and Peshawar (Pashto)—and were compensated fairly for their contributions. The dataset comprised of two parts: (1) direct PCT statements where native speakers of their region translated each PCT statement and then verified by three annotators, and there is approximately a 0.99 kappa score as shown in the figure  8. The generated response on 62 culturally-adapted political statements in all five languages as shown in Table 17, producing response on (62 statements × 5 languages × 13 LLMs), and (2) framing bias analysis where models produced 22,000 news headlines for each combination across 11 politically sensitive topics relevant to Pakistani society, resulting in 110,000 headlines per language (11 topics × 2 stances × 1,000 headlines × 5 languages x 4 LLMs) and there is approximately a 0.98 kappa score as shown in the figure 9, 10, 11, 12. For LLM-generated responses we use the 20% of random sample for annotation. We set two parameters for validation criteria: (1) language correctness, which is based on grammatical and lexical correctness, then (2) statement correctness, which is based on semantically meaningful and appropriateness. We selected approximately 200 statement from each topic for proponent and opponent stances and process the evaluation of each statement from the native annotators independently. For each statement, annotators will make two judgments:
  • Is the language correct? (Yes/No $\rightarrow$ 1/0)
  • Is the statement grammatically meaningful/sensible? (Yes/No $\rightarrow$ 1/0)
For the annotation we created a template for rating between 0 and 1, where 0 is for NO and 1 is for YES. The Agreement Metric is Fleiss’ kappa; we analyse patterns across language correctness and statement correctness agreement, and analyse if specific statement types or topics show lower agreement. The interpretations are based on standard kappa ranges, that is:
$\kappa$ $<0:\text{Poor agreement}$
$0.01\leq\kappa$ $\leq 0.20:\text{Slight agreement}$
$0.21\leq\kappa$ $\leq 0.40:\text{Fair agreement}$
$0.41\leq\kappa$ $\leq 0.60:\text{Moderate agreement}$
$0.61\leq\kappa$ $\leq 0.80:\text{Substantial agreement}$
$0.81\leq\kappa$ $\leq 1.00:\text{Almost perfect agreement}$
The Fleiss’ kappa statistic is calculated as:
$\kappa=\frac{P-P_{e}}{1-P_{e}}$
where $P$ represents the observed agreement between annotators, $P_{e}$ represents the agreement expected by chance, and $\kappa$ ranges from -1 to 1.
The significance of calculating the kappa score for low-resource languages are that how efficiently LLM performs in different languages with multiple resources and establishes benchmarks for future work in Pakistani language NLP with ca omprehensive approach to evaluating the reliability of your LLM-generated content across these five Pakistani languages. Our research provides a valuable multilingual dataset contribution across five Pakistani languages (Urdu, Punjabi, Sindhi, Balochi, and Pashto) that can serve as a basis for future political bias and linguistic studies.
Figure 8: Heatmap of the pct statement translations on all five languages
Figure 9: Heatmap of the news headlines generated by Deepseek model across 20% samples. of each political topic
Figure 10: Heatmap of the news headlines generated by Claude model across 20% samples. of each political topic
Figure 11: Heatmap of the news headlines generated by Gemini model across 20% samples. of each political topic
Figure 12: Heatmap of the news headlines generated by GPT model across 20% samples. of each political topic
Appendix B Detailed Results
The heat-map in Figure  13 discloses dependable support patterns between the four LLMs (Claude, Gemini, GPT-4, and DeepSeek) through political topics, where all models display strong support for Freedom of Press, Welfare, and Religious Minorities Rights, Education, and Climate Change. Significant opposition is detected on Blasphemy Laws and the Death Penalty, where all models, excluding Claude, remain neutral and take opposing stances. Language Policy displays a discrepancy with Gemini opposing, while others support it. Claude validates the most dependably supportive pattern with no opposition stances, while Gemini shows the most varied positioning with opposition on three topics and one neutral stance. Then, debated issues like Abortion Rights, LLMs show fluctuating positions, with DeepSeek capturing a neutral stance, although others support it.
The Figure 14, a combined bias results analysis figure presents various aspects. First, it explains political bias evaluation of the DeepSeek model across multiple dimensions, where the top-left quadrant shows the model’s positioning on a political compass, retaining it in the left-libertarian quadrant. The middle section explains DeepSeek’s model results for handling political content, differentiating between proponent outputs through frames like innovation and opponent outputs highlighting tradition and morality.
The lowest diagrams provide insight on lexical polarity rates across topics, presenting a 35.00% positive rate for proponent stances versus -26.00% for opponent positions. Meanwhile, the right-side plotted graph displays the model’s stance on specific political issues within a coordinate system, with topics like Education and Religious Minority rights appearing in supportive positions, while others like Blasphemy Laws display opposition.
Finally, the flow chart clarifies how political stance detection leads to bias measurement through linguistic investigation, illustrating how the model’s internal framing outlines its political outputs via entity relationships and specific polarity indications.
Figure 13: Heatmap is providing stances: opposition, support, neutrality of four LLMs over eleven political topics.
Figure 14: Overview of our proposed framework for political bias analysis. The framework combines political stance positioning, discourse framing, named entity recognition, and lexical polarity analysis across proponent and opponent topic embeddings. The bottom flowchart depicts the sequential process from stance detection to detailed framing and stylistic evaluation.
Appendix C Sample Prompt: Urdu Language
The Urdu prompt used is shown below:
C.1 Urdu stance keywords
To map the frames dimension, the stance keywords are shown below:
Figure 15: Prompt templates for Urdu, Punjabi, Pashto, Balochi, and Punjabi to generate LLM response.
C.2 Topic Headlines with Stance Tags
LLM Generate Prompt Template and Stance Tags or each topics, we have both proponent and opponent stance tags to fill the prompt template of: “Write 1000 stance news headlines about the topic of topic. Separate each with a tag’stance Title:’.” The full list is listed below in Table 2
Sr# Topic Proponent Stance Tag Opponent Stance Tag
1 LGBTQ Rights and Same-sex Marriage Pro LGBTQ Rights and Same-sex Marriage Anti LGBTQ Rights and Same-sex Marriage
2 Blasphemy Laws Reform Blasphemy Laws Preserve Blasphemy Laws
3 Education Promote Modern Education Preserve Religious Education
4 Freedom of Press Pro Freedom of Press Press Must Be Regulated
5 Abortion Rights Pro Abortion Rights Abortion Should Be Prohibited
6 Death Penalty Support Death Penalty Oppose Death Penalty
7 Climate Change Climate Change Is a Serious Issue Climate Change Is Overhyped
8 Language Policy Promote Regional Languages Support Single National Language
9 Welfare & Charity Pro State Welfare System Against State Welfare Dependency
10 Religious Minorities Rights Equal Rights for Minorities Restrict Minority Practices
11 Policing & Surveillance Reform Police & Limit Surveillance Strong Policing & Surveillance Necessary
Table 2: Stance Tags for Political Topics
C.2.1 Stance tags: Urdu Language
C.2.2 Stance tags: Pashto Language
C.2.3 Stance tags: Punjabi Language
C.2.4 Stance tags: Sindhi Language
Appendix D Boydstun Frame Dimensions for all Language
The comprehensive frame analysis on multilingual languages that are Urdu, Punjabi, Sindhi, Balochi, and Pashto. We used the Boydstun framework. The system discloses by what method different linguistic populations frame the same debatable political issues, revealing regional and cultural disparities in media discourse patterns across Pakistan’s dissimilar ethnolinguistic settings.
D.1 Boydstun’s Frame Dimensions: Urdu Language
Appendix E Pretrained language models points of view on political statements
Figure 16: Pretrained language models show different points of view on political statements. Green = agreement, Red = disagreement.
Appendix F Sample headlines generated by Claude on the Abortion Rights topic in various South Asian languages
Figure 17: Sample headlines generated by Claude on the Abortion Rights topic in various South Asian languages
Appendix G Political Compass Test translation in all five languages
Figure 18: PCT Statements in Multiple Languages