### v2

# Framing Political Bias in Multilingual LLMs Across Pakistani Languages

Abstract
Large Language Models (LLMs) increasingly shape public discourse, yet most evaluations of political and economic bias have focused on high-resource, Western languages and contexts. This leaves critical blind spots in low-resource, multilingual regions such as Pakistan, where linguistic identity is closely tied to political, religious, and regional ideologies. We present a systematic evaluation of political bias in 13 state-of-the-art LLMs across five Pakistani languages: Urdu, Punjabi, Sindhi, Pashto, and Balochi. Our framework integrates a culturally adapted Political Compass Test (PCT) with multi-level framing analysis, capturing both ideological stance (economic/social axes) and stylistic framing (content, tone, emphasis). Prompts are aligned with 11 socio-political themes specific to the Pakistani context. Results show that while LLMs predominantly reflect liberal-left orientations consistent with Western training data, they exhibit more authoritarian framing in regional languages, highlighting language-conditioned ideological modulation. We also identify consistent model-specific bias patterns across languages. These findings show the need for culturally grounded, multilingual bias auditing frameworks in global NLP.
[…]
202F
Framing Political Bias in Multilingual LLMs Across Pakistani Languages
Afrozah Nadeem, Mark Dras, Usman Naseem School of Computing, Macquarie University, Australia, afrozah.nadeem@students.mq.edu.au, {mark.dras,usman.naseem}@mq.edu.au
1 Introduction
Large Language Models (LLMs) have achieved strong performance across a range of NLP tasks and languages Blodgett et al. (2020). However, increasing evidence shows that these models encode social and ideological biases, including hallucinations, stereotypes, and political partisanship Zheng et al. (2023). Political bias is particularly consequential, as it can influence public discourse, reproduce dominant ideologies, and marginalise minority perspectives Demszky et al. (2019).
Bias in language models is not a theoretical concern—it shapes real-world outputs such as news headlines, reinforcing dominant ideologies while marginalizing dissent. This can distort public discourse, erode democratic values, and undermine trust in AI systems Barkhordar et al. (2024). While English and other high-resource languages have received some scrutiny Weidinger et al. (2021), low-resource languages like Urdu, Punjabi, Sindhi, Pashto, and Balochi remain severely underexplored Kumar et al. (2023). In Pakistan, where political identity is closely tied to language, this oversight risks amplifying bias across culturally sensitive issues. Addressing this gap is vital to building fair, inclusive AI systems for linguistically diverse and politically complex societies.
Figure 1: An illustration of political bias in multilingual LLMs using a real Urdu-language response to PCT Statement 24. The model interprets a traditional justice principle as endorsing violence, revealing a liberal bias that overlooks its cultural and religious significance—highlighting ideological misalignment and polarization in Urdu-language LLM responses.
We address this by focusing on Pakistan, a linguistically and politically diverse setting that remains largely absent from current literature as shown in Figure 1, that traditional justice principle as violent, revealing liberal bias and overlooking its cultural and religious significance. Over 80% of its population speaks one of five major languages—Urdu, Punjabi, Pashto, Sindhi, and Balochi—each associated with distinct ideological, ethnic, and religious identities Sun et al. (2022). Political discourse in Pakistan often revolves around highly contested issues such as blasphemy, minority rights, and federalism Harman (2018), making it a valuable test case for analyzing multilingual political bias. Existing approaches to bias evaluation often rely on Western political taxonomies Chen et al. (2020), overlook the framing of ideologies in low-resource languages, and treat languages as isolated units Bang et al. (2021). Moreover, most methods emphasize stance classification while neglecting how style and narrative framing encode bias Yu et al. (2023). Recent work has critiqued the Political Compass Test (PCT) for its prompt sensitivity and lack of cultural grounding in multilingual contexts Röttger et al. (2024), calling for more context-aware evaluations. Our work addresses this by not only adapting the PCT to the sociopolitical landscape of Pakistani languages, but also introducing narrative framing analysis, offering a more nuanced and culturally robust approach to assessing political bias in large language models. We propose a novel framework for evaluating political and economic bias in LLMs across five Pakistani languages. This is the first framework to combine ideological positioning (via PCT) with narrative framing analysis for political discourse in Pakistani languages. This study makes the following contributions:
  • We conduct the first large-scale political bias evaluation in five Pakistani languages.
  • We adapt and translate the PCT to cover 11 culturally salient topics grounded in Pakistani discourse.
  • We propose a three-part framing analysis using Boydstun’s taxonomy, named entity recognition, and lexical polarity.
  • We analyze 13 SOTA LLMs to investigate how political positions and framing strategies vary across languages, and how linguistic choice activates culturally specific ideological shifts.
2 Related Work
2.1 Political Bias in Language Models:
The political orientation of LLMs has emerged as a core concern in AI ethics. Studies show models like GPT-3/4 reflect liberal social leanings and partisan patterns Liu et al. (2021); Motoki et al. (2024); Ceron et al. (2024), but these insights remain western-centric. Tools like the Political Compass Test (PCT) Hartmann et al. (2023) and policy probes Bang et al. (2021) assume linguistic translatability, which fails in low-resource contexts. Bias detection tools often underperform in non-English settings due to cultural misalignment Barkhordar et al. (2024). Pakistani languages—Punjabi, Sindhi, Pashto, and Balochi—are critically underserved, requiring culturally grounded evaluation strategies Harman (2018); Thapa et al. (2024).
2.2 Framing and Discourse-Level Analysis:
Most political bias research emphasizes stance detection, neglecting how bias manifests through rhetorical framing. The taxonomy by Boydstun et al. (2014) provides a foundation for deeper analysis of issue framing, yet remains underused in LLM evaluations. Framing is particularly relevant in multicultural settings, where political language varies not just in content but in style, tone, and structure—areas that remain largely unexplored in multilingual NLP.
Figure 2: Overview of our proposed framework for political bias analysis for evaluating political bias in language models. The framework features a political compass approach for stance detection and decomposes bias into content and style dimensions, examining controversial topics across Pakistani languages.
2.3 Existing Work on Pakistani Languages
Pakistani languages—Urdu, Punjabi, Sindhi, Pashto, and Balochi—are spoken by over 200 million people globally, including large diasporas in the UK, Canada, UAE, and the U.S. Mostefa et al. (2012); Hussain (2004). Despite this, they remain critically under-represented in NLP. Recent work has addressed Urdu QA Arif et al. (2024), data augmentation for NER Ehsan and Solorio (2025), and benchmarking LLMs on Urdu tasks like Sentiment Analysis, Fake News DetectionTahir et al. (2025). However, political bias and framing remain unexplored. Studies highlight how LLMs fail in low-resource contexts due to cultural misalignment Barkhordar et al. (2024); Thapa et al. (2024), reinforcing the need for culturally grounded analysis in South Asian languages.
Despite recent advances, political bias research in language models remains limited by a Western-centric focus and reliance on high-resource languages. Recent work has critiqued the use of the PCT, highlighting its sensitivity to language and lack of cultural grounding Röttger et al. (2024). Addressing this gap, our study presents the first comprehensive framework to evaluate political bias in Pakistani languages, combining a culturally adapted PCT with multi-layered framing analysis. Beyond stance scoring, we incorporate policy frames, named entity recognition, and lexical polarity to capture both ideological alignment and the rhetorical expression of political ideas. We evaluate 13 state-of-the-art LLMs across 11 politically salient topics, providing a unified methodology for analyzing political bias in low-resource, non-Western settings Bang et al. (2024).
3 Method
Overview: Figure 2 presents the comprehensive methodological framework underpinning our political bias analysis. First, we focus on the quantitative evaluation of political stance, utilizing the Political Compass Test to position model outputs across economic and social dimensions. Second, we expand the analysis by examining framing bias, incorporating content framing, named entity recognition, and lexical polarity to capture the nuanced ways models express ideological positions within culturally relevant political topics. Each tier is discussed in detail in the following sections.
3.1 Political Leaning of pretrained LLM
Overview and Motivation: To evaluate political bias in large language models (LLMs), we adopt an evaluation approach grounded in political spectrum theories, following the framework proposed by Feng et al. (2023). This framework characterizes political opinion using a two-axis model: social values, ranging from liberal to conservative ($S_{\mathrm{soc}}$), and economic values, ranging from left to right ($S_{\mathrm{eco}}$). This representation offers a nuanced view of political ideologies, surpassing the simplistic left-right dichotomy and enabling deeper analysis of how political bias manifests in LLMs Liu et al. (2024).
Political Compass Test (PCT): The PCT [^1] https://www.politicalcompass.org/test consists of 62 statements, denoted as $P_{\mathrm{s}}$, designed to assess a respondent’s political orientation across the $S_{\mathrm{eco}}$ and $S_{\mathrm{soc}}$ axes Smith et al. (2022). Each participant evaluates the statements using a four-level Likert scale: Strongly Agree, Agree, Disagree, and Strongly Disagree. The final scores are calculated based on weighted responses and mapped to a coordinate in the range [10,10][-10,10][ - 10 , 10 ] for each axis, producing a two-dimensional representation of political ideology.
Language Models and Prompt Design: We evaluated thirteen LLMs categorized into open-source and closed-source models. The closed-source models include: OpenAI, OpenAI o1-mini, OpenAI o1-preview, GPT-3.5-turbo, GPT-4, GPT-4-turbo, GPT-4o, Claude, and Gemini 1.5 Pro. The open-source models include: Mistral, DeepSeek, RoBERTa-large, RoBERTa-base, and BERT-large, model details are given in Appendix B.3. For each language, we developed tailored prompts to generate high-quality responses, which are shown in Appendix 23.
Stance Scoring and Mapping: To quantify the model’s political stance on each PCT statement, we employ a two-stage computational approach. First, we calculate a scalar stance score from classifier confidence values, for more details see Appendix B.4. Second, we discretize this continuous score into ordinal categories that align with the PCT response format. We discretize stance by selecting the class with the highest softmax confidence score among four candidate labels. The continuous stance score is then calculated by multiplying the confidence value with a fixed weight: $\pm$10 for Strongly Agree/Disagree and $\pm$5 for Agree/Disagree. This approach reflects both the categorical stance and the model’s certainty. This allows us to convert raw classifier outputs into interpretable symbolic judgements. Let $A_{s},A,D,D_{s}\in[0,1]$ denote the predicted confidence scores corresponding to the classes Strongly Agree, Agree, Disagree, and Strongly Disagree, respectively.
1. Stance Score Computation: We define a scoring function $f:[0,1]^{4}\rightarrow[-10,10]$ to assign a value based on the dominant class:
$f(A_{s},A,D,D_{s})=\begin{cases}10\cdot A_{s}&\text{if }A_{s}>\max(A,D,D_{s})\\ 5\cdot A&\text{if }A>\max(A_{s},D,D_{s})\\ -10\cdot D_{s}&\text{if }D_{s}>\max(D,A,A_{s})\\ -5\cdot D&\text{otherwise}\end{cases}$
The result $S=f(A_{s},A,D,D_{s})\in[-10,10]$ serves as a continuous stance score, indicating both the direction and strength of agreement. 2. Stance Discretization: To facilitate comparative analysis across models and statements, we define a discretization function $g:[-10,10]\rightarrow\{0,1,2,3\}$ that maps the continuous score to categorical labels using a symmetric threshold parameter $\tau>0$:
$g(S)=\begin{cases}3&\text{if }S\geq 2\tau\\ 2&\text{if }0\leq S<2\tau\\ 1&\text{if }-2\tau<S<0\\ 0&\text{if }S\leq-2\tau\end{cases}$ (1)
[…]
3.2 Ideological Framing Analysis
While the PCT quantifies political orientation along economic and social dimensions, it lacks detailed insight into how these ideologies are expressed in discourse Rozado (2024). To address this, we propose an ideological framing analysis framework that examines how large language models (LLMs) communicate politically sensitive topics through content and stylistic choices Liu et al. (2024). This method complements PCT by analyzing not only the stance but also the narrative strategies LLMs employ to present their positions Abdurahman et al. (2024).
Topic Selection and Data Generation: We select eleven politically salient topics in the Pakistani context, each mapped to PCT dimensions and characterized by polarized opinions. Topics were identified based on prior research Lee et al. (2022), reputable institutions (e.g., Pew Research Center [^2] https://www.pewresearch.org/topics/), and media bias trackers (e.g., Allsides.com [^3] https://www.allsides.com/topics-issues) Bang et al. (2024). The topics, denoted $T=\{T_{(i)}\}$, include:
LGBTQ Rights and Same-Sex Marriage, Blasphemy Laws, Education, Freedom of Press, Abortion Rights, Death Penalty, Climate Change Ejaz et al. (2023), Language Policy, Welfare and Charity, Religious Minorities Rights, Policing and Surveillance.
Each topic was translated into five Pakistani languages to enable multilingual evaluation. For each topic $T_{(i)}$, we generate news headlines $H_{\mathrm{LLM}}$ in two opposing stances: proponent $T_{(i)\mathrm{pro}}$ and opponent $T_{(i)\mathrm{opp}}$. Headlines are an ideal unit for framing analysis as they encapsulate the core message and tone of discourse Lee et al. (2022); Sheng et al. (2021); Baly et al. (2020). We generated 1000 headlines per stance, per language, using prompts that explicitly specify stance to elicit contrasting viewpoints Nadeem et al. (2021) (see Appendix 23 for prompt details and Appendix B.2for prompting strategy and reproducibility details).
Frame Dimension Classification: To examine ideological narratives in model-generated content, we classify headlines using Boydstun’s 15 cross-cutting frame dimensions (Boydstun et al., 2014), which encompass salient themes such as Economics, Morality, Health and Safety, and Cultural Identity. These topic-independent frames enable consistent comparative analysis across models and topics (Hamborg, 2020). We employ GPT-3.5-turbo with bilingual prompts to classify each headline into one or more frames, enhancing contextual understanding in Pakistani languages. For each topic–stance pair $(t,s)$, we compute the frame ratio for frame $f_{i}$ as:
$\text{FrameRatio}_{t,s}(f_{i})=\frac{c_{t,s}(f_{i})}{N_{t,s}}$ (2)
[…]
Entity-Based Framing Analysis: To explore how models frame specific actors or institutions, we extract named entities—such as political figures, countries, and organizations—from Urdu headlines using a multilingual NER model. For each topic–stance pair $(t,s)$, we count how often each entity $e_{i}$ appears, denoted as $c_{t,s}(e_{i})$. We then calculate its relative Prominence P using:
$\text{P}_{t,s}(e_{i})=\frac{c_{t,s}(e_{i})}{\sum_{j=1}^{n}c_{t,s}(e_{j})}$ (3)
This score reflects which entities are most emphasized in model outputs, offering insights into how narratives center around particular individuals or groups. Frequent entity mentions serve as a subtle framing device, indicating which actors or groups models emphasize Devlin et al. (2019); Zheng et al. (2023). NER was conducted using a pretrained bert-base-multilingual-cased model  [^4] https://huggingface.co/google-bert/bert-base-multilingual-cased.
[…]
$\Pr^{(k)}_{t,s}(e_{i})=\frac{S_{e_{i}}^{(k)}}{\sum_{k^{\prime}}S_{e_{i}}^{(k^{\prime})}}$ (4)
We define the dominant sentiment polarity for entity $e_{i}$ as:
$\text{SentimentBias}_{t,s}(e_{i})=\arg\max_{k}\left(\Pr^{(k)}_{t,s}(e_{i})\right)$ (5)
These scores expose entity-level sentiment bias patterns across topics and stances (Spliethöver et al., 2022; Roy and Goldwasser, 2020).
4 Experimental Settings
Dataset: We introduce and release a novel multilingual dataset designed to support political bias analysis across five Pakistani languages: Urdu, Punjabi, Sindhi, Balochi, and Pashto [^5] https://anonymous.4open.science/r/PoliticalBiasEvaluation-10DE. The dataset comprises two main components:
  • Political Compass Test (PCT) Translations: This segment includes 62 culturally adapted political statements translated into each of the five target languages as detailed Table in Appendix 28. Responses were collected from 13 large language models (LLMs), generating a total of 4,030 responses (62 statements × 13 models × 5 languages). we employed three native speakers per language with triple verification to ensure full linguistic and semantic fidelity. Inter-annotator agreement for these translations achieved a Fleiss’ kappa score of 0.99, indicating near-perfect consensus.
  • Headline Generation Corpus: To evaluate framing bias, we generated news headlines using four SOTA LLMs across 11 politically sensitive topics in both proponent and opponent stances tags see details in Appendix 6. Each model produced 22,000 headlines per language, resulting in a total of approximately 444,340 multilingual headlines (22,000 × 4 models × 5 languages). A stratified 20% sample (2,200 headlines) was manually evaluated for linguistic correctness and semantic coherence, yielding a 100% language correctness rate and a semantic agreement and semantic consistency reflecting moderate inter-annotator agreement given the subjective nature of generated content, as Shown in Figure 3.
The dataset addresses a critical gap in non-Western language resources for political discourse analysis and AI bias evaluation. All annotators were recruited from linguistically representative Urdu– national language, Punjabi – most widely spoken regional language, Sindhi– western provincial language, Balochi – southwestern provincial language,Pashto – northwestern provincial language.—and were compensated fairly. In our dataset, English model responses were collected using the unaltered version of the original PCT. This serves as a baseline for identifying shifts in political stance when the same models are prompted in Pakistani languages.
Figure 3: Kappa score heatmap illustrating Language Correctness (LC) and Statement Correctness (SC) for five regional languages across
[…]
Computational Resources: This study required substantial computational resources to evaluate political bias across multilingual LLMs. We incurred approximately \$287 USD in OpenAI API usage for five languages, alongside cloud expenses for running open-source models and NLP pipelines. The total budget was around \$350 USD. The complete pipeline—including Political Compass Test evaluation, generation of 440,340 headlines, and multi-layer framing analysis—consumed approximately 120 GPU-hours on NVIDIA A100 instances.
5 Results and Analysis
5.1 Political Stance Distribution Across Languages
Figure 4 presents political stance outcomes from the Political Compass Test (PCT) across five Pakistani languages. Most LLMs cluster in the libertarian-left quadrant, reflecting progressive economic and independent social values. Claude shows the strongest libertarian stance, while GPT-4-turbo leans most economically left. Models like GPT-3.5-turbo and OpenAI o1-mini shift toward authoritarian-right in Urdu, highlighting language-specific influences. Sindhi remains consistently libertarian-left, whereas BERT variants lean right across languages. GPT models trend authoritarian-left in regional contexts, unlike open-source models which remain more liberal-libertarian. These findings underscore the need for multilingual political bias evaluation to ensure culturally equitable model behavior Johnson and Goldwasser (2016).
Figure 4: Political leaning of open source and closed source models used for Pakistani language shows diverse inclination across LLM
Cross Language Bias Patterns: Our findings highlight significant cultural variation in LLM behavior. While English outputs tend to align with fairness and neutrality—often in the libertarian-left quadrant—responses in Urdu, Punjabi, and Sindhi shift toward authoritarian-right stances, influenced by cultural norms or training data as shown in Figure 5. Pashto remains closest to English in political leaning. Even fine-tuned Urdu models retain or amplify these biases (Figure 13). Error analysis (Figure 16) shows English as the most stable, whereas Pakistani languages exhibit greater variance and bias. This underscores the need for culturally aware debiasing and targeted evaluation to ensure fairness in multilingual, low-resource LLM applications.
Figure 5: Deterministic decoding reveals political bias shifts in LLMs across six Pakistani languages and English, exposing deep cross-linguistic bias under deterministic conditions.
Figure 6: Entity-level analysis of Urdu language outputs across LLMs. The bar chart (left) shows entity prediction frequency for key institutions, while the right chart visualizes the top 10 entities associated with the “Religious and Minority Rights” topic. with circle sizes indicating mention frequency and colors representing sentiment reflecting entity prominence, highlighting model-specific focus and cultural alignment in politically sensitive contexts.
Cultural Adaptive Models are Less Biased: We have performed additional experiments on the Urdu language to evaluate the political bias on LLMs, to analyse it, we fine-tuned models: Bert-base [^6] https://huggingface.co/google-bert/bert-base-cased,Gemma7b [^7] https://huggingface.co/google/gemma-7b, GPT-3.5, and Roberta-small—all specifically adapted for the Urdu language as shown in Figure 13. The model political inclination can be verified by the bias score of a language model based on its political positioning in a 2D ideological space, which can be measured as $\text{Bias Score}=\sqrt{(x-0)^{2}+(y-0)^{2}}=\sqrt{x^{2}+y^{2}}$. Where: $x$ is the position on the economic scale, $y$ is the position on the social (authoritarian/libertarian) scale, and (0,0)(0,0)( 0 , 0 ) represents perfect neutrality (i.e., bias score = 0). Other bias score formulations are Bias Score = |x| + |y|  [^8] Manhattan Distance or simple magnitude Bias Score = Max(|x|, |y|)  [^9] Chebyshev Distance. The interpretation scale is:
$\textbf{LM}=\left\{\begin{array}[]{ll}\text{Perfect neutrality}&\text{if Bias Score }=0\\ \text{Highly neutral}&\text{if Bias Score }<1\\ \end{array}\right.$
Model classification by bias score is measured as:
$\textbf{LM }=\left\{\begin{array}[]{ll}\text{Neutral model}&\text{if Bias Score }\in[0,1)\\ \text{Biased model}&\text{if Bias Score }\geq 1\\ \end{array}\right.$
Any deviation from the origin (0,0)(0,0)( 0 , 0 ) indicates increasing political bias, while proximity reflects neutrality. Urdu fine-tuned models outperform SOTA baselines by producing more balanced, centered responses. This demonstrates that cultural-to-linguistic adaptation enables context-aware generation, promoting politically neutral and culturally relevant outputs for multilingual AI in diverse regional settings.
5.2 Framing Bias Analysis
Content Bias: In content bias analysis, we evaluated framing dimension and entitiy frequency analysis, which are described below.
Framing Dimension: All models exhibit varying uses of cultural identity frames when discussing religious minority issues in Pakistan as shown in Figure 12. Claude emphasizes universalist fairness, while GPT-4 integrates fairness, morality, and regulation. For the death penalty, Claude and Gemini rely on morality, whereas GPT-4 and DeepSeek favor policy frames. Shared use of constitutional framing reflects Islamic legal influence. Abstract frames lead to higher model error rates, as shown in Figure 7 Fazal (2022).
Entity Frequency Analysis: Entity frequency analysis for Religious Minority Rights reveals framing patterns aligned with political orientations Schramowski et al. (2022). Figure 6 shows that models exhibiting more authoritarian stances tend to frame minority rights through institutional and geographic hierarchies, emphasizing state and regional across top 10 entities, such as Pakistan, the Supreme Court, and regional bodies serve to situate the discourse within legal and geographic contexts, words like rights and law reinforce a rights-based framing. Libertarian-leaning models present a broader spectrum of entities, incorporating both legal frameworks and regional minority experiences, indicating a more nuanced framing.
Stylistic Bias: Lexical polarity analysis highlights how LLMs stylistically frame within Pakistani political discourse. Figure 14 shows some of society’s most sensitive and debated issues. Same-sex marriage rights stand out with the most polarized coverage—yet interestingly, the sentiment leans slightly positive, hinting at a complex and emotionally charged discourse. Language policy, welfare and charity, and education follow closely, marked by passionate debate but generally hopeful tone. On the other hand, deeply rooted religious and moral issues like abortion, blasphemy laws, and the death penalty show intense division and overwhelmingly negative sentiment. When it comes to government performance, coverage tends to be both critical and sharply divided—reflecting growing public frustration as shown in Figure 17. Such stylistic tendencies suggest a diplomatic approach by LLMs to sensitive issues, emphasizing rights and dialogue over conflict or aggression. For more analysis and results see Appendix C.
Figure 7: Error rates across political discourse dimensions reveal how confidently models handle different types of content.
6 Findings
6.1 Model-Level Bias Interpretation
Figure  8 highlights variation in lexical tone, with topics such as Language Policy and education showing consistently high positive LPR, while Blasphemy Laws, Death Penalty, and Religious Minorities Rights exhibit strong negative LPRs, especially for Gemini. The results highlights variation in lexical tone, with topics such as Language Policy and education showing consistently high positive LPR, while Blasphemy Laws, Death Penalty, and Religious Minorities Rights exhibit strong negative LPRs, especially for Gemini. To further interpret how political bias manifests in model behaviour, we conducted a detailed, multidimensional analysis of the DeepSeek model’s visual and quantitative breakdowns, which are provided in the Appendix Figure 21.
Figure 8: Lexical Polarity Rate (LPR) comparison across 11 sociopolitical topics for LLMs.
6.2 Cross-Cultural model insights
Our analysis reveals distinct model behaviors across cultural-linguistic contexts. GPT models show high cultural sensitivity, shifting from libertarian-left in English to authoritarian-left in Pakistani languages (Figure 4). Claude remains consistently libertarian with progressive sentiment. Open-source models vary—Deep Seek offers stable entity selection, while Gemini shows notable ideological shifts. These findings challenge universal bias mitigation, emphasizing that effective global AI requires balancing fairness with cultural nuance—something current models often fail to achieve.
6.3 Multilingual Influence on Political Bias
The clustering plots of political stance detection across languages reveal that LLMs exhibit biased behavior influenced by political and cultural context. This variation highlights imbalances in training data, with English-dominant models adapting differently in low-resource languages. Figure 18 shows consistent support for topics like Education but sharp divergence on Language Policy. These findings suggest that multilingualism can generate language-specific political personalities, raising concerns that global LLM deployment may unintentionally reinforce cultural biases depending on the language used.
7 Conclusion
This study presents the first large-scale investigation of political bias in large language models across five low-resource yet widely spoken Pakistani languages: Urdu, Punjabi, Sindhi, Pashto, and Balochi. We introduce a novel, culturally adapted evaluation framework that integrates ideological stance scoring with discourse-level framing analysis—capturing both what the model says and how it says it. Our findings uncover systematic, language-conditioned shifts in political bias, demonstrating how linguistic and cultural context significantly shapes model behavior. By bridging the Political Compass Test (PCT) with narrative framing dimensions, our methodology offers a reproducible and extensible approach for diagnosing political bias in multilingual LLMs. This work fills a critical gap in multilingual NLP and contributes a practical auditing tool for building culturally grounded, fair, and context-aware AI systems.
Limitations
This study has several methodological and scope-related limitations. Our analysis is restricted to five Pakistani languages, omitting other regional or global languages that may exhibit different ideological patterns. We also exclude high-resource languages like French or Arabic, which could have served as cultural counterfactuals to help distinguish between linguistic and cultural influences in LLM pretraining. While we adapt the Political Compass Test (PCT), its original design is rooted in Western political thought and may not fully capture culturally specific structures like Islamic jurisprudence or tribal governance. Additionally, our sentiment and entity analysis tools are primarily trained on English, potentially reducing accuracy and missing localized expressions. We do not explore prompt-induced framing bias, which may affect stance independently of model ideology. Finally, the static nature of training data limits our ability to assess real-time political shifts. Future work should explore culturally sensitive bias mitigation for low-resource settings.
Ethical Statement
This research was conducted with strict devotion to ethical principles, ensuring cultural sensitivity and participant welfare. Content generation carefully avoided potentially harmful or inflammatory material while maintaining analytical integrity. We acknowledge possible biases in our Western-developed evaluation frameworks and commit to transparent reporting of limitations. The dataset excludes personally identifiable information and extreme political content that could incite violence or discrimination. We recognize the responsibility of AI bias research in multicultural contexts and emphasize that our findings should inform inclusive AI development rather than reinforce stereotypes. This work aims to promote reasonable AI systems that respect diverse political perspectives and cultural values across Global South communities.
Appendix A Appendix
A.1 Dataset Contribution
Our research provides a valuable multilingual dataset that spans five Pakistani languages (Urdu, Punjabi, Sindhi, Balochi, and Pashto) that can serve as a basis for future political bias and linguistic studies. For the Political Compass Test, the statement are translated into five languages shown in the Table 28. To work with five Pakistani languages (Urdu, Pashto, Sindhi, Balochi, and Punjabi), the design study proposed three annotators from the specific region of Pakistan, and each of them are highly expert in speaking and writing in the dedicated low-resource language of their area. The dataset addresses a critical gap in non-Western language resources for political discourse analysis and AI bias evaluation. We chose the Method: Triple-verified by native speakers. All annotators were recruited from linguistically representative regions—Lahore (Urdu), Sahiwal (Punjabi), Karachi (Sindhi), Quetta (Balochi), and Peshawar (Pashto)—and were compensated fairly for their contributions. The dataset comprised of two parts: (1) direct PCT statements where native speakers of their region translated each PCT statement and then verified by three annotators, and there is approximately a 0.99 kappa score as shown in the figure  9. The generated response on 62 culturally adapted political statements in all five languages as shown in Table 28, producing response on (62 statements × 5 languages × 13 LLMs), and (2) framing bias analysis where models produced 22,000 news headlines for each combination across 11 politically sensitive topics relevant to Pakistani society, resulting in 110,000 headlines per language (11 topics × 2 stances × 1,000 headlines × 5 languages x 4 LLMs) and there is approximately a 0.98 kappa score as shown in the figure 3. For LLM-generated responses we use the 20% of random sample for annotation. We set two parameters for validation criteria: (1) language correctness, which is based on grammatical and lexical correctness, then (2) statement correctness, which is based on semantically meaningful and appropriateness. We selected approximately 200 statement from each topic for proponent and opponent stances and process the evaluation of each statement from the native annotators independently. For each statement, annotators will make two judgments:
  • Is the language correct? (Yes/No $\rightarrow$ 1/0)
  • Is the statement grammatically meaningful/sensible? (Yes/No $\rightarrow$ 1/0)
A.2 Language Translation Procedure
For each of the five target languages, prompts were translated from English by a team of three bilingual native speakers. This was followed by model generation in the respective language, forming a two-step pipeline. While multi-translator input reduced individual bias, we acknowledge the absence of back-translation or inter-annotator agreement as a limitation that may introduce subtle framing or cultural interpretation bias in multilingual political contexts.
A.3 Translation Assignment Procedure
Each statement was independently translated into the target language by three bilingual native speakers. Translations were then compared, and a final version was selected through majority agreement or collaborative consensus when needed. This approach ensured semantic accuracy while reducing individual translator bias. For the annotation we created a template for rating between 0 and 1, where 0 is for NO and 1 is for YES. The Agreement Metric is Fleiss’ kappa; we analyse patterns across language correctness and statement correctness agreement, and analyse if specific statement types or topics show lower agreement. The interpretations are based on standard kappa ranges, that is:
[…]
The Fleiss’ kappa statistic is calculated as:
$\kappa=\frac{P-P_{e}}{1-P_{e}}$
where $P$ represents the observed agreement between annotators, $P_{e}$ represents the agreement expected by chance, and $\kappa$ ranges from -1 to 1.
Figure 9: Heatmap of the pct statement translations on all five languages
The significance of calculating the kappa score for Pakistani languages are that how efficiently LLM performs in different languages with multiple resources and establishes benchmarks for future work in Pakistani language NLP with ca omprehensive approach to evaluating the reliability of your LLM-generated content across these five Pakistani languages. Our research provides a valuable multilingual dataset contribution across five Pakistani languages (Urdu, Punjabi, Sindhi, Balochi, and Pashto) that can serve as a basis for future political bias and linguistic studies.
Category Details
PCT
Political statements 62 $P(s)$
Languages 5
LLMs 13
$R_{\text{LLM}}$ 4,030
Agreement score Fleiss’ kappa = 0.99 (near-perfect)
Headline Generation Corpus
Political topics 11 $T(i)$
Stance tags $T(i)_{\text{Pro}}$, $T(i)_{\text{Opp}}$
LLMs 4
Headlines per language per model 22,000
$H_{\text{LLM}}$ 444,340 (approx.)
Human evaluation sample size 2,200 headlines (20% sample)
Language correctness rate 99%
Semantic agreement rate 98.5%
Table 1: Summary Statistics for Political Compass Test Translations and Headline Generation Corpus
Appendix B Technical details
B.1 Stance Detection vs. Our Ideological Mapping
Standard stance detection typically involves classifying whether a text expresses a pro, con, or neutral position toward a specific target or claim Mohammad et al. (2016). These approaches focus on binary or ternary stance concerning an explicit target, often in single-turn texts such as tweets. In contrast, our methodology uses an adapted Political Compass Test to infer a model’s position in a two-dimensional ideological space (economic and social axes). Rather than target-specific classification, we aggregate stance scores across 62 political statements to construct a holistic ideological profile per model and language. This offers a structured lens into political bias beyond isolated stance decisions. Moreover, we complement this scalar stance mapping with rhetorical framing analysis to examine how ideological leanings are expressed stylistically and narratively—going beyond traditional stance detection’s limited focus on polarity or agreement.
Stance detection is a widely studied task in NLP Gorrell et al. (2019), typically formulated as predicting whether a speaker is in favor or against a known target or topic. These tasks are often applied to tweets, debates, or news articles. In contrast, our use of the Political Compass Test allows for continuous stance scoring across a spectrum of ideologically salient statements, enabling two-dimensional mapping of model behavior. This richer representation is particularly important for analyzing latent political bias in generative LLMs, where stance is not tied to a single topic but emerges across diverse ideological domains.
B.2 Prompting Strategy and Reproducibility
We adopt a zero-shot prompting strategy for all LLM generations to avoid bias from exemplars. For each of the 62 PCT statements, models are asked to respond in the specified language using a Likert-style agreement format. For framing analysis, we generate 1,000 headlines per stance–topic pair using bilingual prompts that explicitly request either supportive or opposing tone (see Appendix 21). Frame classification is conducted using GPT-3.5-turbo with a fixed schema based on Boydstun’s taxonomy. Named entity and sentiment analyses are automated via mDeBERTa and XLM-RoBERTa. All code and prompts are shared publicly for reproducibility.
B.3 Model Details
Table 2 provides an overview of the language models used in our bias evaluation. It includes both closed-source and open-source models, detailing their type, estimated parameter sizes, and architectures. Hyperlinked model names direct to official documentation or repositories, enabling transparency and reproducibility for further comparative analysis.
Model Architectures and Tuning Details Most of the models we evaluate, including GPT-4, Claude, Gemini, and DeepSeek, are not raw pretrained models but represent fully developed systems with instruction tuning and safety alignment, often including RLHF. Our focus is on assessing bias as it appears in real-world, user-facing outputs. We include both decoder-based models (e.g., DeepSeek-Chat) and encoder-based models (e.g., BERT, RoBERTa) to explore architectural effects. While decoder models are prompted generatively, encoder models are probed using classification on the same inputs. We acknowledge that this mix, along with post-training layers, may influence results and that further work is needed to isolate these factors.
Table 2: Overview of Language Models Used in Bias Evaluation
Model Name Type Parameters Architecture
GPT-3.5-turbo Closed-source $\sim$175B (est.) Decoder
GPT-4-turbo Closed-source $\sim$1.8T (est.) Decoder
GPT-4 Closed-source $\sim$1.8T (est.) Decoder
GPT-4o Closed-source $\sim$1.8T (est.) Decoder
OpenAI o1-mini Closed-source Unknown Decoder
OpenAI o1-preview Closed-source Unknown Decoder
Claude-3-Haiku-202403 Closed-source $\sim$13B (est.) Decoder
Gemini-1.5-Pro Closed-source Unknown Decoder
Mistral-7B-Instruct-v0.2 Open-source 7B Decoder
DeepSeek-Chat Open-source 7B Decoder
XLM-RoBERTa-large Open-source 550M Encoder
BERT-large Open-source 340M Encoder
XLM-RoBERTa-base Open-source 270M Encoder
B.4 On Bias and Limitations of mDeBERTa
While mDeBERTa-v3-base-mnli-xnli offers strong cross-lingual performance for zero-shot stance classification, we acknowledge that it may carry latent biases inherited from its training on the XNLI corpus. The XNLI dataset is primarily derived from translations of English data and may over-represent high-resource languages and Western discourse norms. This could influence how disagreement or ambiguity is expressed in lower-resource languages like Urdu or Balochi. Although mDeBERTa outperformed alternatives such as XLM-R and mBERT in pilot tests for our target languages, we note that future work should explore culturally fine-tuned models or adversarial probing to better surface language-specific classification bias.
Role of mDeBERTa in Stance Classification. To perform stance classification over multilingual PCT responses, we utilized mDeBERTa-v3-base, a multilingual variant of DeBERTa pretrained on XLM-R corpora, which offers enhanced cross-lingual representation capabilities. We selected mDeBERTa over alternatives such as XLM-RoBERTa and mBERT due to its superior performance in zero-shot stance and sentiment classification benchmarks, especially for underrepresented languages. Its disentangled attention mechanism and language-agnostic pretraining make it a suitable choice for capturing ideological nuance across the five Pakistani languages evaluated. We fine-tuned mDeBERTa on translated PCT examples and constrained the output to four stance labels (Strongly Agree, Agree, Disagree, Strongly Disagree). While mDeBERTa performs competitively, we acknowledge potential limitations from English-centric pretraining that may introduce biases or misalignments in culturally specific contexts, which we mitigate through triple-verified human translations and language-specific prompt tuning. Future work can explore culturally grounded multilingual encoders tailored to South Asian political discourse.
B.5 Stance Classification Details
We use the mDeBERTa-v3-base-mnli-xnli model from HuggingFace’s Transformers library as a zero-shot classifier to assign stance labels. The model is prompted with concatenated input: the PCT statement and model response. It returns softmax confidence scores across four labels: Strongly Agree, Agree, Disagree, and Strongly Disagree. The label with the highest confidence is selected as the predicted stance. We compute a numeric stance score by multiplying the winning label’s score by ±5 or ±10. For example, a confidence of 0.86 on Strongly Disagree yields a stance score of –8.6. This scoring method provides interpretable agreement strength in both ordinal and continuous forms.mDeBERTa-v3-base-mnli-xnli was chosen for stance classification because of its zero-shot multilingual capabilities and language coverage, outperforming alternatives like XLM-R or monolingual classifiers in cross-lingual consistency.
B.6 Standard PCT Automation
We follow the original two-dimensional structure of the Political Compass Test, consisting of the economic (Seco) and social (Ssoc) axes. Rather than replicate its internal scoring algorithm, we simulate user input by mapping model-generated stance scores to the 4-option Likert scale and input them programmatically into the official PCT interface using Selenium. This yields authentic quadrant-level coordinates directly from the source.
B.7 Language Focus and Fine-tuning Scope.
In this study, we did not conduct full model fine-tuning due to resource constraints and instead relied on zero-shot and instruction-tuned responses from existing large language models (LLMs) across multiple languages. For controlled probing in low-resource settings, we strategically focused on Urdu to examine how political and economic bias manifests in culturally grounded contexts. Urdu was selected based on its linguistic richness, wider resource availability, and its status as the mother tongue of Pakistan.
B.8 Framing Setup
We acknowledge that using a single model response per prompt may introduce sampling variance in frame predictions. Incorporating majority voting across multiple generations, or ensembling across different models, could improve the robustness of frame assignment and reduce random variability. We leave this to future work due to API cost constraints.
B.9 Sentiment Classifier
We employ XLM-RoBERTa (XLM-R) for downstream tasks such as multilingual frame classification due to its robust performance across 100 languages, including low-resource ones like Urdu, Punjabi, and Pashto. Trained on CommonCrawl data in a self-supervised manner, XLM-R provides strong cross-lingual generalization, making it well-suited for tasks where labeled data is scarce or unavailable in the target language. Compared to alternatives like mBERT, XLM-R achieves superior results in cross-lingual transfer, particularly for sentence-level classification tasks, while maintaining consistency across diverse scripts. Its architecture also allows effective integration with frame-tagging pipelines in our zero-shot or few-shot evaluations.
Appendix C Detailed Results
C.1 Ideological Consistency of PCT Responses
Figure 10 illustrates the distribution of model responses to Political Compass Test (PCT) prompts projected along the economic (Seco) and social (Ssoc) axes, with bootstrapped 95% confidence intervals shown as translucent error bars. Each dot corresponds to a single prompt, and its position reflects the average ideological stance expressed by the model. The plot reveals a strong diagonal clustering from the lower-left to the upper-right quadrant, indicating a high correlation between the model’s economic and social leanings. Despite variation introduced through multilingual translations, the narrow spread of error bars for most points suggests stable and consistent model behavior. A few prompts with wider intervals reflect ideologically ambiguous or culturally sensitive content. This visualization offers a fine-grained and interpretable view of model ideology, moving beyond discrete stance labels and enabling deeper insight into alignment patterns across ideological dimensions.
C.2 Ideological Leanings of LLMs Across Political Topics
The heat-map in Figure  18 discloses dependable support patterns between the four LLMs (Claude, Gemini, GPT-4, and DeepSeek) through political topics, where all models display strong support for Freedom of Press, Welfare, and Religious Minorities Rights, Education, and Climate Change. Significant opposition is detected on Blasphemy Laws and the Death Penalty, where all models, excluding Claude, remain neutral and take opposing stances. Language Policy displays a discrepancy with Gemini opposing, while others support it. Claude validates the most dependably supportive pattern with no opposition stances, while Gemini shows the most varied positioning with opposition on three topics and one neutral stance. Then, debated issues like Abortion Rights, LLMs show fluctuating positions, with DeepSeek capturing a neutral stance, although others support it.
Figure 10: Mean stance positions on economic (Seco) and social (Ssoc) axes for PCT prompts, with 95% bootstrapped confidence intervals ($n=10$). The diagonal pattern indicates a strong correlation between ideological dimensions.
Figure 11: Stance comparison Religious Minorities Rights
Figure 12: Frame dimension ratio for "Religious minority rights" topic for four models. Overall similar but most variance observed in "Morality", "Fairness and Equality" frames.
Figure 13: Political leaning of four open-source LLMs used for the Fine-tuning on Urdu language: a deep analysis on cultural nuance of political inclination of LLMs
Figure 14: Deep Seek model topic by polarization in Urdu Headlines: Topics ranked by opinion extremity, with color dots indicating positive (green) or negative (red) sentiment direction.
Figure 15: Scatter plot showing how model confidence relates to prediction uncertainty. When models are highly confident, their bootstrap error can remain significant. The trend line ($y=1.15x-0.19$) reveals a slight increase in error with confidence, suggesting that confidence alone is not a reliable indicator of trustworthiness in multilingual outputs.
C.3 Boydstun Framing Patterns in Pro vs Opp Across LLMs
Figure 22 presents a comparative analysis of framing dimensions across key sociopolitical topics using four large language models: Claude, DeepSeek, Gemini, and GPT-4. Each subplot corresponds to a specific topic (e.g., LGBTQ rights, climate change, education) and illustrates the distribution of Boydstun framing dimensions (e.g., Economic, Morality, Fairness and Equality, Security and Defence). The solid bars represent the percentage of responses invoking each frame, with distinctions made between proponent and opponent stances. Clear patterns emerge: “Morality” dominates discussions on blasphemy laws, “Fairness and Equality” is prominent in LGBTQ-related topics, and “Capacity and Resources” frequently appears in education debates. These results underscore how framing choices vary not only by model, but also by issue and stance—revealing nuanced ideological tendencies embedded in LLM outputs.
C.4 Boydstun Framing Model Confidence
Figure 15 illustrates the relationship between model confidence and the intensity of Boydstun Framing in responses across political discourse. Each point represents a model output, plotted by its confidence score (X-axis) and normalized framing intensity (Y-axis), with color indicating the dominance of a particular frame dimension (Frame Percentage). The fitted regression line (dashed red) reveals a positive linear trend ($y=1.15x-0.19$), suggesting that responses with higher model confidence tend to exhibit stronger or more consistent use of specific frames. This indicates a potential coupling between linguistic certainty and ideological framing, where confident outputs are more likely to reinforce a particular narrative or interpretive lens. The trend supports the hypothesis that framing is not incidental, but may become more pronounced when the model generates responses it deems more certain—highlighting the need for deeper scrutiny in high-confidence predictions when auditing bias in LLM outputs.
C.5 Lexical Polarity
The Figure 20 compares sentiment patterns between proponents and opponents of religious minorities’ rights, considering a topic specifically related to Religious Minority Rights, as shown in Figure 17. It also shows predominantly positive sentiment, while opponents display more polarized views. This indicates that LLMs frame religious minority rights using constructive and humanitarian language, consistent with international human rights norms, while recognizing Pakistan’s complex religious landscape.
This finding highlights a critical limitation in relying solely on confidence scores as indicators of reliability, especially in multilingual settings. The presence of significant bootstrap variance even at high confidence levels underscores the need for more robust uncertainty-aware evaluation frameworks.
C.6 Polarization
Figure 19 shows that claude demonstrates the most positive lexical framing across topics, while Gemini shows the least. In terms of polarization, Gemini exhibits the highest variability in sentiment across topics, whereas GPT-4 maintains the consistent and balanced tone, with the lowest average polarization score.
The results reveal distinct sentiment and polarization profiles across four language models.
Claude exhibits the highest overall average polarity (+0.080), followed by GPT-4 (+0.070) and DeepSeek (+0.060), indicating generally positive framing, while Gemini produces the least positive responses (+0.030). In contrast, Gemini shows the highest average polarization (0.630), suggesting greater variability and potentially more divisive language across topics. DeepSeek (0.610) and Claude (0.550) also display moderate polarization, whereas GPT-4 maintains the lowest polarization (0.510), reflecting more balanced sentiment. These results highlight GPT-4 as the most tonally consistent model, while Claude is the most positive and Gemini the most polarizing.
These results highlight GPT-4 as the most tonally consistent model, while Claude is the most positive and Gemini the most polarizing.
C.7 Political Bias Insights Through Our Framework
The Figure 21, a combined bias results analysis figure presents various aspects. First, it explains political bias evaluation of the DeepSeek model across multiple dimensions, where the top-left quadrant shows the model’s positioning on a political compass, retaining it in the left-libertarian quadrant. The middle section explains DeepSeek’s model results for handling political content, differentiating between proponent outputs through frames like innovation and opponent outputs highlighting tradition and morality.
This includes political compass positioning, stance-specific framing differences, sentiment polarity rates, and topic-wise alignment patterns. The analysis confirms that DeepSeek consistently occupies a left-libertarian space while framing proponent content more positively and emphasizing tradition in opponent discourse. A sequential flowchart outlines the model’s internal decision process across stance, framing, and polarity layers.
The lowest diagrams provide insight on lexical polarity rates across topics, presenting a 35.00% positive rate for proponent stances versus -26.00% for opponent positions. Meanwhile, the right-side plotted graph displays the model’s stance on specific political issues within a coordinate system, with topics like education and religious minorities appearing in supportive positions, while others like blasphemy laws display opposition, as shown in Figure 11.
Finally, the flowchart illustrates how political stance detection leads to bias measurement through linguistic investigation, showing how the model’s internal framing outlines its political outputs via entity relationships and specific polarity indications.
C.8 Do Models from the Same Family Exhibit the Same Bias?
While models within the same architectural family often share foundational characteristics and pretraining objectives, our results reveal that political and framing biases are not strictly consistent across family lines—particularly in multilingual settings. For instance, OpenAI’s GPT series (GPT-3.5, GPT-4, GPT-4o) generally aligns with libertarian-left positions in English but exhibits divergent quadrant shifts in Pakistani languages, such as GPT-3.5 adopting a more authoritarian-right stance in Urdu. Similarly, although Claude models consistently favor fairness-based frames and exhibit ideological stability across languages, Gemini models show pronounced shifts toward legalistic or conservative frames in religious and social topics. These findings suggest that language context, fine-tuning procedures, and task framing significantly mediate the expression of bias, even within the same family. Consequently, model family lineage alone cannot reliably predict ideological behavior—highlighting the need for language-specific and context-aware evaluations of LLM fairness.
C.9 Does Model Size Correlate with Political Neutrality?
Our analysis suggests that while larger language models (e.g., GPT-4, Claude) tend to produce more consistent and coherent responses across multilingual prompts, they are not inherently more politically neutral. For instance, GPT-4 consistently leans libertarian-left in English and retains relatively stable ideological positions across languages compared to its smaller counterpart GPT-3.5, which exhibits greater quadrant drift in regional languages. This indicates improved alignment and consistency with scale. However, neutrality is not guaranteed—larger models may still amplify sociopolitical patterns embedded in their training data. Prior studies Liu et al. (2021); Hartmann et al. (2023) also observe that larger models often internalize liberal-leaning priors due to dominant trends in web-scale corpora. Consequently, model size improves reliability but does not equate to fairness or ideological balance, especially in cross-cultural contexts.
C.10 Framing and Political Bias in NLP.
Political bias in NLP has traditionally been examined through the lens of stance detection or ideological scoring, where models are evaluated based on their alignment with predefined political axes (e.g., liberal vs conservative, authoritarian vs libertarian). However, recent research highlights that how a model expresses its position—its framing—can reveal deeper, subtler forms of bias that are not captured by stance alone. Framing refers to the rhetorical, stylistic, and narrative strategies used to present an issue Boydstun et al. (2014), including which aspects are emphasized, which entities are foregrounded, and what tone is adopted. In politically sensitive domains, framing choices can reinforce dominant ideologies, obscure alternative perspectives, or sway audience interpretation without explicitly stating a stance. Therefore, framing analysis complements traditional stance-based evaluations by uncovering the narrative structure of bias. This is particularly crucial in multilingual and culturally diverse contexts, where political expression varies not only in content but also in stylistic conventions. Our work advances this direction by systematically quantifying framing bias across languages and models, using frame taxonomy, entity prominence, and sentiment polarity as multidimensional indicators of narrative bias.
C.11 Liberal-Leaning Tendencies in Political Topics.
Across multiple evaluations, we observe that state-of-the-art language models tend to exhibit a consistent liberal or left-leaning bias when responding to political prompts—particularly in English and high-resource settings. This trend manifests in both stance scoring and in the framing of sensitive topics such as LGBTQ rights, abortion, welfare, and climate change. For instance, models like GPT-4 and Claude frequently emphasize frames of fairness, equality, and moral responsibility, while minimizing authoritarian or traditionalist perspectives. Such patterns align with prior studies Hartmann et al. (2023), which attribute these leanings to the influence of Western liberal norms embedded in web-scale training data. While alignment tuning may reinforce these biases for safety and inclusivity, it also raises concerns about the ideological neutrality of LLMs—especially when deployed in culturally diverse or conservative regions. Our findings confirm that liberal-leaning responses are not isolated artifacts but rather systemic tendencies that persist across models and languages, albeit modulated by linguistic context and prompt framing.
Table 3: Bootstrap-based Bias Evaluation Metrics for Political Compass Responses Across Models and Languages.
Model Language Avg Conf Min Conf Max Conf Low Conf Pred Bootstrap Error 95% CI Conf-weighted Error Weighted Mean
GPT-4-Turbo English 0.602 0.276 0.960 26/62 (6.5%) $\pm$0.518 [1.304, 2.339] $\pm$2.518 2.567
Urdu 0.658 0.317 0.978 12/62 (19.4%) $\pm$0.786 [-0.480, 1.092] $\pm$3.630 0.188
Pashto 0.575 0.297 0.977 23/62 (37.1%) $\pm$0.735 [-0.704, 0.765] $\pm$3.363 -0.042
Punjabi 0.584 0.284 0.972 21/62 (33.9%) $\pm$0.588 [0.896, 2.072] $\pm$3.066 1.537
Balochi 0.579 0.353 0.889 16/62 (25.8%) $\pm$0.668 [0.176, 1.512] $\pm$2.954 0.877
Sindhi 0.579 0.314 0.931 19/62 (30.6%) $\pm$0.670 [-0.266, 1.075] $\pm$3.191 0.390
Gemini-1.5-Pro English 0.836 0.577 0.975 0/62 (0.0%) $\pm$0.898 [-2.394, -0.598] $\pm$4.015 -1.331
Urdu 0.578 0.317 0.978 25/62 (40.3%) $\pm$0.649 [-2.144, -0.845] $\pm$2.803 -1.748
Pashto 0.560 0.290 0.985 29/62 (46.8%) $\pm$0.607 [-1.679, -0.466] $\pm$2.977 -1.407
Punjabi 0.520 0.275 0.916 38/62 (61.3%) $\pm$0.766 [-1.557, -0.025] $\pm$2.903 -1.287
Balochi 0.551 0.323 0.910 24/62 (38.7%) $\pm$0.633 [-1.757, -0.491] $\pm$2.657 -1.445
Sindhi 0.547 0.260 0.981 32/62 (51.6%) $\pm$0.550 [-2.046, -0.946] $\pm$2.628 -2.170
Mistral-7B English 0.635 0.270 0.934 13/62 (21.0%) $\pm$0.660 [-0.194, 1.125] $\pm$3.664 0.654
Urdu 0.595 0.297 0.956 24/62 (38.7%) $\pm$0.771 [-0.048, 1.494] $\pm$3.320 0.809
Pashto 0.514 0.299 0.946 34/62 (54.8%) $\pm$0.718 [-0.718, 0.718] $\pm$3.020 0.098
Punjabi 0.566 0.337 0.911 24/62 (38.7%) $\pm$0.695 [-1.741, -0.352] $\pm$3.241 -0.892
Balochi 0.592 0.335 0.915 17/62 (27.4%) $\pm$0.729 [-0.878, 0.580] $\pm$3.204 -0.206
Sindhi 0.518 0.328 0.959 36/62 (58.1%) $\pm$0.554 [-1.077, 0.031] $\pm$2.937 -0.607
DeepSeek-Chat English 0.836 0.577 0.975 0/62 (0.0%) $\pm$0.866 [-2.309, -0.577] $\pm$4.015 -1.331
Urdu 0.673 0.298 0.988 13/62 (21.0%) $\pm$0.852 [-1.187, 0.516] $\pm$3.664 -0.542
Pashto 0.628 0.321 0.962 18/62 (29.0%) $\pm$0.830 [-0.252, 1.408] $\pm$3.712 0.451
Punjabi 0.584 0.297 0.934 15/62 (24.2%) $\pm$0.769 [0.499, 2.036] $\pm$3.036 1.383
Balochi 0.579 0.353 0.889 16/62 (25.8%) $\pm$0.668 [0.176, 1.512] $\pm$2.954 0.877
Sindhi 0.631 0.339 0.978 18/62 (29.0%) $\pm$0.838 [-0.555, 1.121] $\pm$3.559 0.145
Table 4: Political Compass Scores Across Languages and Models
Model Name Urdu Punjabi Pashto Sindhi Balochi
Econ. Soc. Econ. Soc. Econ. Soc. Econ. Soc. Econ. Soc.
Closed Source Models
GPT-3.5-turbo 0.5 -0.1 1.38 1.95 -0.13 2.1 1.0 1.49 1.38 1.03
GPT-4-turbo -2.38 -1.54 -2.13 -0.21 -1.63 0.26 1.13 -0.05 2.88 0.97
GPT-4 -1.0 -1.23 1.75 1.08 0.13 2.0 2.38 0.97 1.25 2.1
GPT-4o -1.75 -1.03 -1.5 -2.26 -1.13 -0.97 0.13 -1.03 2.38 1.08
OpenAI o1-mini 0.75 -0.82 0.0 -1.23 1.13 -0.56 1.63 -0.31 -0.13 -0.21
OpenAI o1-preview -1.13 -0.92 1.38 0.31 -1.38 0.51 0.75 0.36 1.5 -0.62
Claude-3-Haiku-202403 0.25 -1.79 1.13 0.15 -2.63 -0.26 0.0 0.72 -1.0 1.59
Open Source Models
Gemini-1.5-Pro -0.75 -2.1 -1.0 0.31 -0.13 -1.03 -0.25 -1.33 1.75 0.77
Mistral-7B-Instruct-v0.2 2.5 1.23 -1.0 0.31 0.0 -0.41 -0.75 -2.26 1.5 1.23
DeepSeek-Chat -1.0 -1.23 -0.25 -0.05 -1.0 0.87 0.38 -1.28 -2.13 1.64
XLM-RoBERTa-large 1.5 0.31 2.38 -0.15 2.0 -0.62 1.75 -0.51 -0.13 1.69
BERT-large 0.5 -0.62 1.63 -0.46 2.0 -0.51 1.75 -0.56 0.0 1.28
XLM-RoBERTa-base 1.88 -0.21 1.63 0.31 1.38 -0.41 1.63 -0.21 1.75 0.97
Table 5: Political Compass Scores for Multilingual and English Responses Across Four LLMs
Model Name Urdu Punjabi Pashto Sindhi Balochi English
Econ. Soc. Econ. Soc. Econ. Soc. Econ. Soc. Econ. Soc. Econ. Soc.
gpt4-turbo -0.25 -1.18 0.00 0.51 -0.25 -1.13 1.00 -0.56 -2.38 -0.46 0.75 -0.62
gemini-1.5-pro -0.75 -2.92 -1.00 -0.87 0.38 -2.00 1.75 -2.31 -0.50 -0.62 -2.25 -0.62
Mistral-7B-Instruct-v0.2 0.38 0.62 -0.63 -1.18 -1.50 -1.95 1.88 -0.97 0.25 0.56 -1.63 -1.18
DeepSeek-chat 0.00 -0.46 0.50 2.28 -0.25 -0.21 0.75 0.00 0.50 0.26 -2.75 -0.97
Figure 16: Error distribution of political stance predictions across LLMs for English and Pakistani languages.
Figure 17: Polarization and Sentiment Trends in Urdu Media Headlines Across Sociopolitical Topics.
Figure 18: Heatmap is providing stances: opposition, support, neutrality of four LLMs over eleven political topics.
Figure 19: Overall average polarity and polarization by model.
Figure 20: Sentiment comparison between proponents and opponents of religious minority rights.
Figure 21: Overview of our proposed framework for political bias analysis. The framework combines political stance positioning, discourse framing, named entity recognition, and lexical polarity analysis across proponent and opponent topic embeddings. The bottom flowchart depicts the sequential process from stance detection to detailed framing and stylistic evaluation.
Figure 22: Framing Dimension Distribution Across Models and Topics
Appendix D Sample Prompt: Urdu Language
The Urdu prompt used is shown below:
Large Language Models (LLMs) are becoming powerful voices in shaping public opinion, yet their political and economic biases in low-resource, non-Western languages remain largely overlooked. In this study, we examine 13 popular LLMs across five Pakistani languages—Urdu, Punjabi, Sindhi, Balochi, and Pashto—using a culturally adapted Political Compass Test and a layered framing analysis. While most models show liberal-left tendencies in English, their responses shift toward more authoritarian tones in regional languages. These shifts reflect cultural influences on model behavior and highlight the need for more inclusive, culturally aware approaches to evaluating and improving AI fairness across languages.
D.1 Urdu stance keywords
To map the frames dimension, the stance keywords are shown below:
Figure 23: Prompt templates for Urdu, Punjabi, Pashto, Balochi, and Punjabi to generate LLM response.
D.2 Topic Headlines with Stance Tags
LLM Generate Prompt Template and Stance Tags or each topics, we have both proponent and opponent stance tags to fill the prompt template of: “Write 1000 stance news headlines about the topic of topic. Separate each with a tag’stance Title:’.” The full list is listed below in Table 6
D.2.1 Stance tags: Urdu Language
D.2.2 Stance tags: Urdu Language
Figure 24: Illustration of Urdu stance tags with bordered image
D.2.3 Stance tags: Pashto Language
D.2.4 Stance tags: Punjabi Language
D.2.5 Stance tags: Sindhi Language
Sr# Topic Proponent Stance Tag Opponent Stance Tag
1 LGBTQ Rights and Same-sex Marriage Pro LGBTQ Rights and Same-sex Marriage Anti LGBTQ Rights and Same-sex Marriage
2 Blasphemy Laws Reform Blasphemy Laws Preserve Blasphemy Laws
3 Education Promote Modern Education Preserve Religious Education
4 Freedom of Press Pro Freedom of Press Press Must Be Regulated
5 Abortion Rights Pro Abortion Rights Abortion Should Be Prohibited
6 Death Penalty Support Death Penalty Oppose Death Penalty
7 Climate Change Climate Change Is a Serious Issue Climate Change Is Overhyped
8 Language Policy Promote Regional Languages Support Single National Language
9 Welfare & Charity Pro State Welfare System Against State Welfare Dependency
10 Religious Minorities Rights Equal Rights for Minorities Restrict Minority Practices
11 Policing & Surveillance Reform Police & Limit Surveillance Strong Policing & Surveillance Necessary
Table 6: Stance Tags for Political Topics
Appendix E Boydstun Frame Dimensions for all Language
The comprehensive frame analysis on multilingual languages that are Urdu, Punjabi, Sindhi, Balochi, and Pashto. We used the Boydstun framework. The system discloses by what method different linguistic populations frame the same debatable political issues, revealing regional and cultural disparities in media discourse patterns across Pakistan’s dissimilar ethnolinguistic settings.
E.1 Boydstun’s Frame Dimensions: Urdu Language
Figure 25: Boydstun’s Frame Dimensions: Urdu Language
Appendix F Pretrained language models points of view on political statements
Figure 26: Pretrained language models show different points of view on political statements. Green = agreement, Red = disagreement.
Appendix G Sample headlines generated by Claude on the Abortion Rights topic in various South Asian languages
Figure 27: Sample headlines generated by Claude on the Abortion Rights topic in various South Asian languages
Appendix H Political Compass Test translation in all five languages
Figure 28: PCT Statements in Multiple Languages